A 0.2kg is drag 0.5m upward with a uniform velocity along a plane inclined 30° with the horizontal by a force parallel to the incline. The coefficient of kinetic friction between the object and the plane is 0.25. How much work is done by the applied force? Show complete solution.

1
Expert's answer
2021-11-30T17:44:37-0500

So, we have "h=0.5\ (m)" , "m=0.2\ (kg)" , "\alpha=30°" , "\mu=0.25" .


"W=\Delta E"


"W_F-W_{Fr}=\Delta E\to W_F=\Delta E+W_{Fr}=mgh+\mu mg\cos\alpha\cdot(h/\sin\alpha)="


"=0.2\cdot9.8\cdot0.5+0.25\cdot0.2\cdot9.8\cdot\cos30°\cdot(0.5/\sin30°)=1.4\ (J)" . Answer